# A Fully Automated Pipeline for Conversational Discourse Annotation: Tree Scheme Generation and Labeling with Large Language Models

### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have shown promise in automating discourse annotation for conversations. While manually designing tree annotation schemes significantly improves annotation quality for humans and models, their creation remains time-consuming and requires expert knowledge. We propose a fully automated pipeline that uses LLMs to construct such schemes and perform annotation. We evaluate our approach on speech functions (SFs) and the Switchboard-DAMSL (SWBD-DAMSL) taxonomies. Our experiments compare various design choices, and we show that frequency-guided decision trees, paired with an advanced LLM for annotation, can outperform previously manually designed trees and even match or surpass human annotators while significantly reducing the time required for annotation. We release all code and resultant schemes and annotations to facilitate future research on discourse annotation.
@@ 1 Introduction
Discourse analysis is essential in NLP tasks like dialog management, generation, summarization, and emotion recognition (Liang et al., 2020; Chen et al., 2021; Shou et al., 2022). Traditionally, discourse annotation depends on manual expert labeling, which is costly and time-consuming. LLM-based annotation presents a promising alternative, enhancing speed, consistency, and cost-effectiveness (Gilardi et al., 2023; Hao et al., 2024). However, challenges such as biases and domain limitations necessitate careful prompt design and continuous evaluation.
@@
In Ostyakova et al. (2023), the authors explored using ChatGPT to automate discourse annotation for labeling chit-chat dialogs using the speech functions (SFs) taxonomy (Eggins and Slade, 2004). SFs categorize communicative acts in dialog, capturing speaker intentions and interactions in a hierarchical structure to analyze conversational flow (see Figure 1 for an example and Appendix A for the full label set). Ostyakova et al. (2023) conducted three sets of experiments: (1) Direct Annotation: The LLM assigned labels from a predefined list of SFs; (2) Step-by-Step Scheme with Intermediate Labels: LLM selected labels progressively from broad to specific categories; (3) Complex Tree-Like Scheme with Yes/No Questions: using a complex tree-like annotation pipeline originally designed for crowdsourced annotation. Since breaking a multi-label selection task into smaller sub-tasks using a tree structure has improved human performance in complex discourse annotation (Scholman et al., 2016), the authors hypothesized that the same approach could enhance LLM-based annotation. Prior research also suggests that guiding models with tree-structured prompts significantly improves performance (Yao et al., 2024).
Ostyakova et al. (2023) found that the Tree-Like Scheme approach enhances LLM accuracy, achieving near-human performance, and suggest that LLMs could serve as a “silver standard” for annotation. However, newer and more powerful LLMs have been released since this work was published. These developments not only enhance the potential of LLMs to be used for annotation but also open the door to automating the creation of tree-like schemes, [^1] We will refer to them as tree schemes. enabling a fully automated pipeline. This is especially valuable for large taxonomies, such as Intelligent PAL (Morrison et al., 2014) in the educational domain and TEACh-DA (Gella et al., 2022) for task-oriented systems, where manually creating tree schemes is complex and time-intensive.
This work automates tree schemes creation, making them usable for annotation by crowdsourced workers and LLMs. A tree scheme is a decision tree that classifies dialog utterances through a series of questions, which can be binary or non-binary, using yes/no or open-ended formats. An example of a tree generated using the pipeline proposed in this work can be found in Appendix B. We provide a repository with the code and generated trees to support further research and applications. [^2] https://github.com/Kpetyxova/autoTree
@@ 2 Related Work
Researchers analyze discourse structures to improve dialog understanding and management, focusing on pragmatics and speaker intent. Two key frameworks are the Dialog Act (DA) Theory Jurafsky et al. (1998), which assigns pragmatic labels to utterances, and Segmented Discourse Representation Theory (SDRT) Lascarides and Asher (2007), which models relationships between discourse units. [^3] These frameworks are presented as general background and are not used in the experiments.
@@
Ostyakova et al. (2023) introduce a semi-automated approach for annotating open-domain dialogs using a taxonomy of speech functions and ChatGPT. Their study evaluates three methods: Direct Annotation (selecting from a complete label set), Step-by-Step (progressively narrowing choices), and Tree-Like Schemes (using hierarchical yes/no questions). The Tree-Like Scheme has performed best, particularly for rare classes, achieving high consistency when running the same annotation pipeline with ChatGPT three times (Fleiss’ kappa: 0.83). However, expert input has remained essential for designing the annotation pipelines.
Yadav et al. (2024) and Tseng et al. (2024) explore GPT-4-based semantic annotation, emphasizing prompt design and model limitations. Chen et al. (2024) investigate LLMs for event extraction, addressing data scarcity in fine-tuned models. Finally, Wu et al. (2024) introduce a rationale-driven collaborative framework that refines annotations through iterative reasoning, outperforming standard prompting.
@@ 3 Speech Functions Corpus
The tag set covers five functional dimensions: turn management, topic organization, feedback, communicative acts, and pragmatic purposes. While all dimensions are embedded within speech functions, they are distributed unevenly across tags, with individual speech functions incorporating between two and five dimensions. Figure 1 illustrates an example of a speech function that includes all five dimensions: React.Rejoinder.Support.Track.Clarify. In this example: (1) React represents turn management, indicating a speaker change or a reaction to a previous utterance; (2) Rejoinder corresponds to topic organization, signifying active topic development that influences the dialog flow; (3) Support denotes feedback, showing that the speaker is supporting an interlocutor; (4) Track falls under communicative acts, identifying questions; (5) Clarify serves a pragmatic purpose, indicating a question aimed at obtaining additional information on the current conversation topic. Some SFs, however, cover fewer dimensions. For instance, Open.Attend represents only two: turn management (marking the conversation’s beginning) and communicative acts (a greeting).
@@ 4 Framework for Tree Construction with LLMs
The pipeline for the tree construction process is illustrated in Figure 2. The core concept of the proposed algorithm is to use an advanced LLM to identify distinguishing features that can divide classes into two or more groups. The LLM is provided with a table containing class names, definitions, and usage examples as input.
Figure 2: Pipeline for tree construction. LLM formulates a classification question to split classes into groups, mapping possible answers to respective class groups. This process repeats recursively for created groups until all groups contain only one class. Finally, the grouped data is merged into a single tree structure in JSON format for annotation.
@@
5 Experiments & Results
5.1 Yes/No Questions vs. Open-Ended Questions
The manually created tree in Ostyakova et al. (2023) is binary and composed exclusively of yes/no questions. In this subsection, we implement a binary yes/no question tree using the proposed framework. Additionally, we construct a binary tree with open-ended questions. The hypothesis behind this experiment is that allowing the model to generate any questions, rather than limiting it to yes/no questions, can result in a more flexible and nuanced tree for discourse annotation. The prompts used for constructing both the yes/no questions and the open-ended questions tree were developed through iterative refinement. This refinement process involved adjusting and testing prompts until they produced stable and consistent results, ensuring a valid JSON output with all necessary keys present. The final versions of these prompts are provided in Appendix I and Appendix J, respectively.
Evaluation results for the Top Level, the Top Two Levels, and Top-20 class subsets on the development set are presented in Tables 12 and 3 while the results for the complete SF taxonomy on the development and test sets are shown in Tables 4 and 5, respectively. [^4] The best results for GPT-3.5 are highlighted in bold, while the second-best results are marked with an underline. The overall best results across all models are highlighted in blue. The findings indicate that open-ended trees outperform yes/no trees across all metrics and data subsets. However, for the complete SF taxonomy, both yes/no and open-ended trees perform significantly worse than the manually created tree.
@@ 5 Experiments & Results
5.2 Binary vs. Non-Binary Trees
Allowing the model to split nodes into multiple groups, rather than restricting it to two groups, has the potential to create a more granular and detailed annotation tree. However, this added granularity may introduce greater complexity and reduce consistency in the annotation process. [^5] This experiment is distinct from the previous one, where we only allowed binary open-ended questions and compared them to yes/no questions. As a result, both the questions and answers in that setting were more high-level, distinguishing data based on a single characteristic. In contrast, non-binary open-ended questions allow for greater granularity and specificity, enabling more nuanced differentiation within the data. Experiments described in this section examine the impact of restricting the model to binary splits compared to allowing multiple-group splits on the accuracy of the resulting annotation tree. Both the binary and non-binary trees discussed in this section use open-ended questions. A prompt that allows open-ended questions and splitting data into more than two groups is detailed in Appendix K.
The findings indicate that allowing the model to split data into multiple groups generally outperforms restricting it to binary splits (Tables 1, 2, 3, 4 and 5). This performance difference is more pronounced when the number of classes is smaller. However, as the number of classes increases, the gap narrows, with weighted metrics occasionally favoring the binary approach, specifically showing (1) higher $P_{w}$ and $R_{w}$ for the Top Two Levels (6 classes); (2) higher $R_{w}$ for the Top-20 classes; (3) higher $P_{w}$ for the complete taxonomy (33 classes) on the test set. Nevertheless, higher macro metrics for non-binary trees suggest improved performance for smaller and less frequent classes. Based on these results, further experiments will allow the model to split data into more than two groups.
5.3 Optimal Split Selection and Backtracking
For the Two-Level subset, performance improves compared to the approach that does not use split selection and backtracking (Table 2). In the Top-20 subset, only the $F1$ metric shows an increase, indicating better performance for less frequent classes (Table 3). For the complete SF taxonomy, performance on the development set remains comparable to the approach without split selection and backtracking (Table 4), while on the test set, the $F1$ score is higher (Table 5). [^7] Results for the Top Level subset are unavailable, as these trees have only one level.
@@ 5.3 Optimal Split Selection and Backtracking
5.4 Frequency-Guided Optimal Split Selection and Backtracking
@@
Manual analysis revealed that during the annotation step, the model frequently selected incorrect paths, often defaulting to upper-level classes. This behavior was especially prevalent in the most frequent class, Sustain.Continue.Prolong.Extend. Despite this, the tree structure appears logical, and the root question is straightforward: Does the dialog utterance provide supplementary or contradictory information to the previous statement by the same speaker? The response options for this question are: (1) Dialog utterances that provide supplementary or contradictory information to the previous statement by the same speaker; (2) Dialog utterances that do not provide supplementary or contradictory information to the previous statement by the same speaker. While the question is specific, emphasizing conditions about the same speaker and the addition of information, the model often ignored these requirements. In numerous cases, the first response option was incorrectly assigned, even at the start of a conversation.
@@
The distinguishing characteristic of this tree is the heightened granularity and specificity of the questions and labels, with each step designed to determine whether the utterance fits a particular class using a single, targeted question. To assess whether the frequency-guided tree presents too significant a challenge for the GPT-3.5 model and whether it might perform better with a more advanced model, GPT-4o Hurst et al. (2024) was used during the annotation step (see Section 5.4.1). This approach allowed for a direct comparison between GPT-4o and GPT-3.5. Section 5.4.2 also examines whether the observed differences in metrics are the same for trees created without frequency guidance.
5.4.1 GPT-4o for Annotation
@@
Evaluation results comparing annotations by GPT-4o and GPT-3.5 on trees constructed using frequency-guided optimal split selection and backtracking for the Two-Level and Top-20 class subsets are presented in Tables 2 and 3. Results for the complete SF taxonomy on the development and test sets are shown in Tables 4 and 5, respectively. We note that not only are the differences in metrics highly significant, but also the $\mathbf{P_{w}}$, $\mathbf{R_{w}}$, and $\mathbf{F1}$ scores for annotations on both the development and test sets for the complete SF taxonomy surpass those obtained for the entire dataset annotated by crowdsourcers in Ostyakova et al. (2023). This finding underscores that the proposed Frequency-Guided Optimal Split Selection approach, combined with an advanced LLM for annotation, may both outperform manually constructed trees and improve traditional human-driven annotation processes.
5.4.2 Do Annotation Gaps Persist in Non-Frequency-Guided Approaches?
@@
This section examines whether the significant differences observed between annotations generated by GPT-4o and GPT-3.5 also occur in the non-frequency-guided optimal split selection and backtracking approach.
@@
Tables 234 and 5 presents performance metrics for annotations produced using GPT-3.5 and GPT-4o with the optimal split selection algorithm, both with and without frequency guidance. For the frequency-guided approach, the performance gap between GPT-3.5 and GPT-4o becomes more pronounced as the number of classes increases. In contrast, there is no significant difference in performance for the non-frequency-guided approach when switching from GPT-3.5 to GPT-4o. This underscores the clear advantage of the frequency-guided approach when paired with a more advanced model.

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have shown promise in automating discourse annotation for conversations. While manually designing tree annotation schemes significantly improves annotation quality for humans and models, their creation remains time-consuming and requires expert knowledge. We propose a fully automated pipeline that uses LLMs to construct such schemes and perform annotation. We evaluate our approach on speech functions (SFs) and the Switchboard-DAMSL (SWBD-DAMSL) taxonomies. Our experiments compare various design choices, and we show that frequency-guided decision trees, paired with an advanced LLM for annotation, can outperform previously manually designed trees and even match or surpass human annotators while significantly reducing the time required for annotation. We release all code and resultant schemes and annotations to facilitate future research on discourse annotation: https://github.com/Kpetyxova/autoTree.
@@ 1 Introduction
Discourse analysis is essential in NLP tasks like dialog management, generation, summarization, and emotion recognition (Liang et al., 2020; Chen et al., 2021; Shou et al., 2022). Traditionally, discourse annotation depends on manual expert labeling, which is costly and time-consuming. LLM-based annotation presents a promising alternative, enhancing speed, consistency, and cost-effectiveness (Gilardi et al., 2023; Hao et al., 2024). However, challenges such as biases and domain limitations necessitate careful prompt design and evaluation.
@@
In Ostyakova et al. (2023), the authors explored using ChatGPT to automate discourse annotation for labeling chit-chat dialogs using the speech functions (SFs) taxonomy (Eggins and Slade, 2004). SFs categorize communicative acts in dialog, capturing speaker intentions and interactions in a hierarchical structure to analyze conversational flow (see Figure 1 for an example and Appendix A for the full label set). Ostyakova et al. (2023) conducted three sets of experiments: (1) Direct Annotation with an LLM assigning labels from a predefined list of SFs; (2) Step-by-Step Scheme with Intermediate Labels with an LLM selecting labels progressively from broad to specific categories; and (3) Complex Tree-Like Scheme with Yes/No Questions, using a complex tree-like annotation pipeline originally designed for crowdsourced annotation. Since breaking a multi-label selection task into smaller sub-tasks using a tree structure has improved human performance in complex discourse annotation (Scholman et al., 2016), the authors hypothesized that the same approach could enhance LLM-based annotation. Prior research also suggests that guiding models with tree-structured prompts significantly improves performance (Yao et al., 2024).
Ostyakova et al. (2023) found that the Tree-Like Scheme approach enhances LLM accuracy, achieving near-human performance, and suggested that LLMs could serve as a “silver standard” for annotation. However, newer and more powerful LLMs have been released since this work was published. These developments not only enhance the potential of LLMs to be used for annotation but also open the door to automating the creation of tree-like schemes, [^1] We will refer to them as tree schemes. enabling a fully automated pipeline. This is especially valuable for large taxonomies, such as Intelligent PAL (Morrison et al., 2014) or TEACh-DA (Gella et al., 2022) for task-oriented systems, where manually creating tree schemes is complex and time-intensive.
This work automates tree schemes creation, making them usable for annotation by crowdsourced workers and LLMs. A tree scheme is a decision tree that classifies dialog utterances through a series of questions, which can be binary or non-binary, using yes/no or open-ended formats. An example of a tree generated using the pipeline proposed in this work can be found in Appendix B.
@@ 2 Related Work
Researchers analyze discourse structures to improve dialog understanding and management, focusing on pragmatics and speaker intent. One of the key frameworks is the Dialog Act (DA) Theory Jurafsky et al. (1998), which assigns pragmatic labels to utterances.
@@
Ostyakova et al. (2023) introduce a semi-automated approach for annotating open-domain dialogs using a taxonomy of speech functions and ChatGPT. Their study evaluates three methods: Direct Annotation (selecting from a complete label set), Step-by-Step (progressively narrowing choices), and Tree-Like Schemes (using hierarchical yes/no questions). The Tree-Like Scheme has performed best, particularly for rare classes, achieving high consistency when running the same annotation pipeline with ChatGPT three times (with Fleiss’ kappa of 0.83). However, expert input has remained essential for designing the annotation pipelines.
In addition, Yadav et al. (2024) and Tseng et al. (2024) explore GPT-4-based semantic annotation, emphasizing prompt design and model limitations. Chen et al. (2024) investigate LLMs for event extraction, addressing data scarcity in fine-tuned models. Finally, Wu et al. (2024) introduce a rationale-driven collaborative framework that refines annotations through iterative reasoning, outperforming standard prompting.
@@ 3 Speech Functions Corpus
The tag set covers five functional dimensions: turn management, topic organization, feedback, communicative acts, and pragmatic purposes. While all dimensions are embedded within speech functions, they are distributed unevenly across tags, with individual speech functions incorporating between two and five dimensions. Figure 1 illustrates an example of a speech function that includes all dimensions: React.Rejoinder.Support.Track.Clarify. In this example: (1) React represents turn management, indicating a speaker change or a reaction to a previous utterance; (2) Rejoinder corresponds to topic organization, signifying active topic development that influences the dialog flow; (3) Support denotes feedback, showing that the speaker is supporting an interlocutor; (4) Track falls under communicative acts, identifying questions; (5) Clarify serves a pragmatic purpose, indicating a question aimed at obtaining additional information on the current conversation topic. Some SFs, however, cover fewer dimensions. For instance, Open.Attend represents only two: turn management (marking the conversation’s beginning) and communicative acts (a greeting).
@@ 4 Framework for Tree Construction with LLMs
The pipeline for the tree construction process is illustrated in Figure 2. The core concept of the proposed algorithm is to use an advanced LLM to identify distinguishing features that allow it to divide a set of classes into two or more groups. Here, a group refers to a subset of the input classes that the model clusters together based on a shared property it identifies – framed through a classification question. For example, given the classes Open.Demand.Fact (requesting factual information at the beginning of a conversation), Open.Demand.Opinion (requesting an opinion at the beginning), Open.Give.Fact (providing factual information at the beginning), and Open.Give.Opinion (providing an opinion at the beginning), the model might divide them as group 1: Open.Demand.Fact, Open.Demand.Opinion, and group 2: Open.Give.Fact, Open.Give.Opinion. To do this, the LLM is provided with a table containing class names, definitions, and usage examples as input.
Figure 2: Pipeline for tree construction. An LLM formulates a classification question to split classes into groups, mapping possible answers to respective class groups. This process repeats recursively for created groups until all groups contain only one class. Finally, the grouped data is merged into a single tree structure in JSON format for annotation.
@@
5 Tree Construction Approaches
Yes/No Questions vs. Open-Ended Questions
The manually created tree in Ostyakova et al. (2023) is binary and composed exclusively of yes/no questions. As the first approach explored in our work, we implement a binary yes/no question tree using the proposed framework. Additionally, we construct a binary tree with open-ended questions. The hypothesis behind this experiment is that allowing the model to generate any questions, rather than limiting it to yes/no questions, can result in a more flexible and nuanced tree for discourse annotation. The prompts used for constructing both the yes/no questions and the open-ended questions tree were developed through iterative refinement. This refinement process involved adjusting and testing prompts until they produced stable and consistent results, ensuring a valid JSON output with all necessary keys present. The final versions of these prompts are provided in Appendix I and Appendix J, respectively.
Binary vs. Non-Binary Trees
Allowing the model to split nodes into multiple groups, rather than restricting it to two groups, has the potential to create a more granular and detailed annotation tree. However, this added granularity may introduce greater complexity and reduce consistency in the annotation process. This experiment is distinct from the previous one, where we only allowed binary open-ended questions and compared them to yes/no questions. As a result, both the questions and answers in that setting were more high-level, distinguishing data based on a single characteristic. In contrast, non-binary open-ended questions allow for greater granularity and specificity, enabling more nuanced differentiation within the data. Experiments described in this section examine the impact of restricting the model to binary splits compared to allowing multiple-group splits on the accuracy of the resulting annotation tree. Both the binary and non-binary trees discussed in this section use open-ended questions. A prompt that allows open-ended questions and splitting data into more than two groups is detailed in Appendix K.
Optimal Split Selection and Backtracking
Frequency-Guided Optimal Split Selection and Backtracking
6 Results & Analysis
Evaluation results for the Top Level, the Top Two Levels, and Top-20 class subsets on the development set are presented in Tables 12 and 3 while the results for the complete SF taxonomy on the development and test sets are shown in Tables 4 and 5, respectively. [^3] The best results for GPT-3.5 are highlighted in bold, while the second-best results are marked with an underline. The overall best results across all models are highlighted in blue.
@@ 6 Results & Analysis
Yes/No Questions vs. Open-Ended Questions
The findings indicate that open-ended trees outperform yes/no trees across all metrics and data subsets. However, for the complete SF taxonomy, both yes/no and open-ended trees perform significantly worse than the manually created tree.
Binary vs. Non-Binary Trees
The findings show that allowing the model to split data into multiple groups generally outperforms restricting it to binary splits. This performance difference is more pronounced when the number of classes is smaller. However, as the number of classes increases, the gap narrows, with weighted metrics occasionally favoring the binary approach, specifically showing (1) higher $P_{w}$ and $R_{w}$ for the Top Two Levels (6 classes); (2) higher $R_{w}$ for the Top-20 classes; (3) higher $P_{w}$ for the complete taxonomy (33 classes) on the test set. Nevertheless, higher macro metrics for non-binary trees suggest improved performance for smaller and less frequent classes. Based on these results, further experiments will allow the model to split data into more than two groups.
Optimal Split Selection and Backtracking
For the Two-Level subset, performance improves compared to the approach that does not use split selection and backtracking. In the Top-20 subset, only the $F1$ metric shows an increase, indicating better performance for less frequent classes. For the complete SF taxonomy, performance on the development set remains comparable to the approach without split selection and backtracking, while on the test set, the $F1$ score is higher. [^4] Results for the Top Level subset are unavailable, as these trees have only one level.
@@
Frequency-Guided Optimal Split Selection and Backtracking
@@
Manual analysis revealed that during the annotation step, the model frequently selected incorrect paths, often defaulting to upper-level classes. This behavior was especially prevalent in the most frequent class, Sustain.Continue.Prolong.Extend. Despite this, the tree structure appears logical, and the root question is straightforward: Does the dialog utterance provide supplementary or contradictory information to the previous statement by the same speaker? The response options for this question are: (1) “Dialog utterances that provide supplementary or contradictory information to the previous statement by the same speaker"; (2) “Dialog utterances that do not provide supplementary or contradictory information to the previous statement by the same speaker". While the question is specific, emphasizing conditions about the same speaker and the addition of information, the model often ignored these requirements. In numerous cases, the first response option was incorrectly assigned, even at the start of a conversation.
@@
The distinguishing characteristic of this tree is the heightened granularity and specificity of the questions and labels, with each step designed to determine whether the utterance fits a particular class using a single, targeted question. To assess whether the frequency-guided tree presents too significant a challenge for the GPT-3.5 model and whether it might perform better with a more advanced model, GPT-4o Hurst et al. (2024) was used during the annotation step (see Section 6.1). This approach allowed for a direct comparison between GPT-4o and GPT-3.5. Section 6.2 also examines whether the observed differences in metrics are the same for trees created without frequency guidance.
6.1 GPT-4o for Annotation
@@
Evaluation results comparing annotations by GPT-4o and GPT-3.5 on trees constructed using frequency-guided optimal split selection and backtracking for the Two-Level and Top-20 class subsets are presented in Tables 2 and 3, while Tables 4 and 5 show the results for the complete SF taxonomy on the development and test sets, respectively. We note that not only are the differences in metrics highly pronounced, but also the $\mathbf{P_{w}}$, $\mathbf{R_{w}}$, and $\mathbf{F1}$ scores for annotations on both the development and test sets for the complete SF taxonomy surpass those obtained for the entire dataset annotated by crowdsourcers in Ostyakova et al. (2023). This finding underscores that the proposed Frequency-Guided Optimal Split Selection approach, combined with an advanced LLM for annotation, may both outperform manually constructed trees and improve traditional human-driven annotation processes.
6.2 Do Annotation Gaps Persist in Non-Frequency-Guided Approaches?
@@
This section examines whether the substantial differences observed between annotations generated by GPT-4o and GPT-3.5 also occur in the non-frequency-guided optimal split selection and backtracking approach.
@@
Tables 234 and 5 presents performance metrics for annotations produced using GPT-3.5 and GPT-4o with the optimal split selection algorithm, both with and without frequency guidance. For the frequency-guided approach, the performance gap between GPT-3.5 and GPT-4o becomes more pronounced as the number of classes increases. In contrast, there is no substantial difference in performance for the non-frequency-guided approach when switching from GPT-3.5 to GPT-4o. This underscores the clear advantage of the frequency-guided approach when paired with a more advanced model.
6.3 Error Distribution across Depth Levels
Table 6 decomposes accuracy by the depth at which an annotator model first diverges from the gold path. Several clear trends emerge. First, the use of frequency-guided optimal split selection and backtracking in combination with GPT-4o for annotation nearly eliminates root-level mistakes: the annotator model correctly selects the branch in 98% of the cases. A similar trend is evident at the second level of the tree, where using GPT-4o with a frequency-guided tree yields 95% accuracy, significantly outperforming the 68% accuracy observed with GPT-3.5 in the same context.
Split-Sel. Freq-Guided
Level Metric GPT-3.5 GPT-4o GPT-3.5 GPT-4o
1 Acc. 94.71 74.07 85.71 98.41
Err21.28 55.06 25.00 6.38
2 Acc. 93.85 82.14 67.74 95.12
Err23.40 28.09 46.30 17.02
3 Acc. 87.35 86.73 78.10 80.13
Err44.68 16.85 21.30 65.96
4 Acc. 97.18 100.00 72.00 98.48
Err4.26 0.00 6.48 2.13
5 Acc. 86.36 100.00 88.89 88.24
Err6.38 0.00 0.93 8.51
6 Acc. - - 100.00 100.00
Err- - 0.00 0.00
Table 6: Accuracy and error percentages (relative to the total number of errors) at each tree depth level produced by GPT-3.5 and GPT-4o, using trees constructed with frequency-guided and non-frequency-guided split selection algorithms.
For GPT-4o on frequency-guided trees, errors primarily propagate downward, concentrating at the third level of the tree, where 66% of all errors occur. Most third-level errors arise from misclassifying React.Rejoinder.Support.Response.Resolve (response that provides the information requested in the question) as React.Respond.Confront.Reply.Disagree (negative response). These misclassifications occur because the annotator model confuses the group label “Utterances that involve a positive or negative response to a previous statement” with “Utterances that provide the information requested in the question.” The semantic overlap between these groups indicates that the issue lies more in the class definitions rather than a serious annotation error. Another frequent source of error involves the vague Other utterances group. Utterances of class React.Respond.Support.Register (e.g., “Yeah,” “Hmm…,” “Right”) were grouped under Other utterances during tree construction but later misannotated as React.Respond.Support.Reply.Affirm (positive answers or confirmations, e.g., “Yes”). Despite explicit instructions to avoid creating Other groups during the tree-creation stage, the model occasionally disregarded these instructions and included them, contributing to misclassification. Notably, fewer than 10% of errors occur beyond the third level, suggesting most challenges arise earlier in the tree.
6.4 Consistency of Tree Generation
An important aspect to evaluate is the consistency of the proposed framework: if the decision tree is generated multiple times, how similar or different will the resulting trees be? Table 7 shows annotation performance on the test set using three trees built with the frequency-guided split selection algorithm. The first and third runs yield nearly identical results, while the second performs slightly worse. Manual inspection confirms that all three trees are broadly similar, with the drop in the second run due to its poor handling of the frequent class – React.Respond.Support.Develop.Extend. The primary difference between this second tree and the others lies in its structure: it immediately separates the labels Extend, Enhance, and Elaborate into three distinct terminal nodes at a single step. In contrast, the other two trees first group Enhance and Elaborate together, distinguishing them from Extend, and only in the subsequent step split the remaining group into separate nodes. Overall, the framework demonstrates strong consistency across runs.
Run $\mathbf{P_{w}}$ $\mathbf{R_{w}}$ $\mathbf{F1_{w}}$ $\mathbf{F1}$
1 0.77 0.68 0.67 0.46
2 0.66 0.64 0.62 0.41
3 0.76 0.66 0.68 0.43
Table 7: Evaluation of annotations on the test set using trees generated across three runs with the frequency-guided optimal split selection approach. Annotations were produced by GPT-4o.